# Head-wise Shareable Attention for Large Language Models

### v1

@@ 5.2.2 Evaluation on PostShare
Last, it is important to emphasize that here we perform post-training with limited training corpus and thus it runs the risk of overfitting when training only for one epoch. For example, PostShare achieves the higher accuracy in BoolQ at 0.3 epoch than at 0.5 epoch (68.29 vs. 66.57). In contrast, as the training epoch increases from 0.5 to 0.9, the accuracy in WinoGrande rises (52.98 vs. 54.39). It means that due to the domain-constrained corpus, overfitting to one specific dataset will potentially compromise the capabilities in other tasks. The in-depth analysis is provided in Appendix D.1.
@@ 5.3 Additional Analysis
Table 5: Weight Sharing and Quantization on Llama 2-7B.
Ablation on Impact of Different Head-wise Matching Functions. For weight sharing, the choice of shared heads is critical. In Figure 3, we plot the performance curve on PIQA (Bisk et al., 2019) and OpenBookQA using different head-wise match functions for Baichuan 2-7B model. And the corresponding detailed results are presented in Appendix D.2. Notably, using the cosine similarity between the concatenation matrix of $W^{q}$ and $W^{k}$ attains the most favorable outcomes. This may be because it guarantees the maximum similarities between attention maps from the model before and after weight sharing. Also, this choice is much more stable and robust in some tasks like reading comprehension(e.g., OpenBookQA).
Robustness on the Model Size. In previous experiments, we adopt our approach in LLMs settings. Since small-scale models are not highly over-parameterized as large-scale models (Gao et al., 2023), we further verify the effectiveness of our method on smaller models like BERT-base, GPT2-small. For this analysis, we set the sharing ratio from 0% to 50% with a step of 10% for the fine-tuned GPT-small model on WMT-14 En-Fr dataset. As shown in Table 6, at a 50% sharing ratio, the GPT-small can still yield a BLEU score of 39.44 without any post-training. Such kind of variance in performance is acceptable that to some degree proves our method is also suitable for small-scale models.
@@
Combine Weight Sharing with Quantization. In terms of saving memory consumption, post-quantization employs the strategy of reducing precision in the LLM parameters, while weight sharing aims to reduce the number of parameters. From these two different directions, we suppose integrating weight sharing and quantization may help towards even more memory reduction of LLMs. Hence, we choose GPTQ (Frantar et al., 2022) as a representative and test the effectiveness of applying two techniques in tandem. Specifically, we quantize Llama 2-7B model after 30% DirectShare to 4 bit precision. As is reported in Table 5, they can be effectively combined with no more than 5 points performance drop.
@@
Figure 5: Results in Various Benchmarks via DirectShare+PostShare based on Llama 2-7B model.
Visualization Study on the Shared Weights. To provide a more detailed explanation of our rationale behind head-wise weight sharing, we conduct a visualization study on the ratios of weight sharing across the MHA layers in two models of different scales (see Appendix D.3). Results indicate the shareable weights distribution across attention heads is similar regardless of the sharing ratio. We also observe a relative balanced sharing ratio across MHA layers than layer-wise weight sharing, which may seem counter-intuitive. However, we find such fine-grained operation on weights has already been used in model pruning Sun et al. (2023); Ma et al. (2023), constantly superior to layer-wise pruning.
@@ 6 Conclusion
In this paper, we illustrate the feasibility of fine-grained weight sharing strategy applied in LLMs, namely, head-wise shareable attention. Consequently, we propose two methods for head-wise weight sharing called DirectShare and PostShare, which are complementary in terms of time and performance. Our DirectShare concentrates on a simple, no-training yet effective sharing strategy, performing competitively with one of the state-of-the-art model pruning methods. PostShare, on the other hand, shows an impressive performance on keeping LLM’s capabilities, needing to compromise on time efficiency. Last, we hope our work inspires researchers to explore better fine-grained weight sharing techniques for memory-efficient LLMs.
@@
This paper primarily focuses on the head-wise weight sharing in Multi-Head Attention (MHA) block, inspired by the attention map similarity across heads. However, the Feed-Forward Network (FFN) block has more parameters compared to the MHA block. To further reduce the memory usage for LLMs, there is necessary to investigate the feasibility of applying weight sharing to FFN block. Subsequently, similar to exploration in MHA block, we should determine whether layer-wise weight sharing in FFN block is enough, otherwise fine-grained shared modules are needed to keep more performance. We leave it as future work.
@@
Appendix A Implementation Details
@@
A.1 For the Baseline
To our knowledge, there is no existing baseline for our methods, due to the absence of prior research on fine-grained weight sharing for LLMs. To provide a comprehensive demonstration of the effectiveness of our DirectShare, we can only choose another important memory-efficient method of a different category for comparison. Here, we select two model pruning methods applied in LLMs: one classical model pruning method Magnitude Pruning and one state-of-the-art structured pruning method LLM-Pruner. We do not consider unstructured pruning methods in this paper since they can not achieve real memory reduction without specialized hardware or software.
@@ A.1 For the Baseline
A.2 For the Post-training
For carrying out the post-training process, we employ the code framework from LLaMA-Factory repository [^1] https://github.com/hiyouga/LLaMA-Factory with DeepSpeed ZeRO-1222Because of our designed special loss function in the post-training stage, only DeepSpeed ZeRO-1 can work.. The Adam optimizer with a learning rate of 5e-5 is used in our experiment and the parameter values assigned during training are $\beta_{1}=0.9$ and $\beta_{2}=0.95$. For Llama 2-7B model, we set the batch size to 32. While for Llama 2-13B model, the batch size of training is only 8 subject to the limited computational resources. Besides, the maximum context size and $\gamma$ are set to 4096 and 0.5, respectively.
Appendix B Experimental Results based on Baichuan 2 Models
@@
B.1 Logical and Common Sense Reasoning
Table 7 presents a comparison on five datasets about reasoning abilities for three memory-efficient methods performed on the Baichuan 2 models.
Our results show that compared to NLU and knowledge-related abilities (listed in Table 8,9), DirectShare can indeed maintain its reasoning abilities to a large extent. Specifically, at 30% ratio, DirectShare remains competitive with LLM-Pruner.
@@ B.1 Logical and Common Sense Reasoning
Table 7: Evaluation Results on Reasoning of the Memory-efficient Baichuan 2-7B & Baichuan 2-13B.
B.2 Natural Language Understanding
@@
Table 8: NLU Abilities of the Memory-efficient Models.
B.3 Knowledge-related Tasks
The results of Baichuan 2 models on knowledge-related tasks are shown in Table 9. Similar decline appears in Llama 2-7B and Llama 2-13B models as well.
@@ B.3 Knowledge-related Tasks
Table 9: Results on Knowledge-related Tasks of the Memory-efficient Models.
Appendix C PostShare on Llama 2-13B Model
@@
0% Llama 2-13B 55.44 71.50 40.17 55.81 60.24 58.03 42.40 60.00
30% DirectShare 50.18 59.36 22.30 30.79 26.53 27.53 27.40 27.80
$*{}^{*}$ 53.68 $\uparrow 3.50$ 71.25 $\uparrow 11.89$ 25.80 $\uparrow 3.50$ 33.90 $\uparrow 3.11$ 32.03 $\uparrow 3.30$ 29.07 $\uparrow 1.54$ 33.60 $\uparrow 6.20$ 38.80 $\uparrow 11.00$
Table 10: Performance of the Memory-efficient Llama 2-13B via PostShare. *** means choosing relatively good performance across different training steps.
Appendix D More Analysis
(a) Sharing Ratio=20%
(b) Sharing Ratio=30%
Figure 6: Ratios of Weight Sharing across the MHA Layers in Llama 2-7B/13B & Baichuan 2-7B/13B.
D.1 Overfitting Phenomenon in PostShare
@@
D.2 Impact of Different Head-wise Matching Functions
@@
D.3 Visualization Study on the Shared Weights

### v2

@@ 5.2.2 Evaluation on PostShare
Last, it is important to emphasize that here we perform post-training with limited training corpus and thus it runs the risk of overfitting when training only for one epoch. For example, PostShare achieves the higher accuracy in BoolQ at 0.3 epoch than at 0.5 epoch (68.29 vs. 66.57). In contrast, as the training epoch increases from 0.5 to 0.9, the accuracy in WinoGrande rises (52.98 vs. 54.39). It means that due to the domain-constrained corpus, overfitting to one specific dataset will potentially compromise the capabilities in other tasks. The in-depth analysis is provided in Appendix F.2.
@@ 5.3 Additional Analysis
Statistics of Memory Reduction. Table 3 presents the statistics of the parameter count and memory requirements when applying DirectShare. When sharing 30% parameter sharing in the MHA block, our method achieves 10-13% memory.
Moreover, we find our weight sharing strategy (Section 4.1) also applies to FFN block. We directly observe the weight matrix similarity in FFN and find the concatenation matrix of gate_proj, up_proj and down_proj can be used as matching function for FFN block. Since FFN does not have explainable fine-grained sub-blocks (like attention heads in MHA), we use Traversal Searching method to choose the optimal size of sub-block and find sharing the whole FNN layer works best in the performance maintenance. Finally, when we share 30% of parameter sharing in both MHA and FFN block, the model can save 26-28% GPU memory.
Sharing Ratio #Params GPU Memory
Llama2-7B
0% 6.74B/100% 17826M/100%
30% MHA 6.09B/90.36% 15512M/87.02%
30% MHA+FFN 4.74B/70.33% 12932M/72.55%
Llama2-13B
0% 13.02B/100% 30800M/100%
30% MHA 11.76B/90.32% 27898M/90.58%
30% MHA+FFN 9.21B/70.74% 23002M/74.68%
Table 3: The actual memory savings brought by DirectShare on Llama2 models (recorded during inference on the BoolQ Dataset in OpenCompassv1.0 platform).
@@
Table 4: Performance of combining weight sharing and quantization on Llama2-7B model.
Ablation on Head-wise Matching Functions. For weight sharing, the choice of shared heads is critical. In Figure 3, we plot the performance curve on PIQA (Bisk et al., 2020) and OpenBookQA using different head-wise match functions for Baichuan2-7B model. And the corresponding detailed results are presented in Appendix F.3. Notably, using the cosine similarity between the concatenation matrix of $W^{q}$ and $W^{k}$ attains the most favorable outcomes. This may be because it guarantees the maximum similarities between attention maps from the model before and after weight sharing. Also, this choice is much more stable and robust in some tasks like reading comprehension(e.g., OpenBookQA).
Robustness on the Model Size. In previous experiments, we adopt our approach in LLMs. Since small-scale models are not highly over-parameterized as large-scale models (Gao et al., 2023), we further verify the effectiveness of our method on smaller models like BERT-base, GPT2-small. For analysis, we set the sharing ratio from 0% to 50% with a step of 10% for the fine-tuned GPT-small model on WMT-14 En-Fr dataset. As shown in Table 5, at a 50% sharing ratio, the GPT-small can still yield a BLEU score of 39.44 without any post-training. Such kind of variance in performance is acceptable that to some degree proves our method is also suitable for small-scale models.
@@
Combine Weight Sharing with Quantization. In terms of saving memory, post-quantization employs the strategy of reducing precision in the LLM parameters, while weight sharing aims to reduce the number of parameters. From these two different directions, we suppose integrating weight sharing and quantization may help towards even more memory reduction of LLMs. Hence, we choose GPTQ (Frantar et al., 2022) as a representative and test the effectiveness of applying two techniques in tandem. Specifically, we quantize Llama2-7B model after 30% DirectShare to 4 bit precision. As reported in Table 4, they can be effectively combined with no more than 5 points performance drop.
@@
Figure 5: Evaluation results when combining DirectShare with PostShare based on Llama2-7B model.
Visualization Study on the Shared Weights. To provide a more detailed explanation of our rationale behind head-wise weight sharing, we conduct a visualization study on the ratios of weight sharing across the MHA layers in two models of different scales (see Appendix F.6). Results indicate the shareable weights distribution across attention heads is similar regardless of the sharing ratio. We also observe a relative balanced sharing ratio across MHA layers than layer-wise weight sharing, which may seem counter-intuitive. However, we find such fine-grained operation on weights has already been used in model pruning Sun et al. (2023); Ma et al. (2023), constantly superior to layer-wise pruning.
@@ 6 Conclusion
In this paper, we illustrate the feasibility of fine-grained weight sharing strategy applied in LLMs, namely, head-wise shareable attention. Consequently, we propose two methods for head-wise weight sharing called DirectShare and PostShare, which are complementary in terms of time and performance. Our DirectShare concentrates on a simple, no-training yet effective sharing strategy, performing competitively with one of the state-of-the-art model pruning methods. PostShare, on the other hand, shows an impressive performance on keeping LLM’s capabilities, needing to compromise on time efficiency. Last, we hope our work inspires researchers to explore better fine-grained weight sharing techniques for LLMs.
@@
This paper primarily focuses on the head-wise weight sharing in Multi-Head Attention (MHA) block, inspired by the attention map similarity across heads. Although we have explored the feasibility of our proposal weight sharing strategy in the Feed-Forward Network (FFN) block, we only complete downstream evaluation on Baichuan2-7B model. To further verify the effectiveness of applying weight sharing to both MHA and FFN block, we should offer comprehensive experimental validation across different models and compare the results with baselines. We leave it as future work.
@@
Appendix A Detailed Explanation for DirectShare
Algorithm 1 summarizes the procedure of DirectShare:
(1) Matching (line 1-13)
We first initialize the buffer $\mathcal{D_{\tau}}$ (line 1) which is then used to store each candidate shareable attention head pairs. Next, the iterative process of matching begins:
Prepare Candidate Pairs (line 2-9): Given the number of layers $\mathcal{L}$ and the number of heads per MHA block $\mathcal{H}$, we can construct candidate attention head pairs for sharing. For each pair, we then record the layer index and the head index accordingly, i.e., $index_{i}$ and $index_{m}$ respectively.
Calculate Similarity (line 10-13): We use 10 randomly selected samples from Wikipedia as the calibration samples for calculating the weight matrix similarity. We present the design of head-wise match function in Section 4.1. The final matching score $\mathcal{S}_{cos}$ is obtained by averaging across samples (line 10). For every attention head, our approach selects the most matching counterpart according to $\mathcal{S}_{cos}$ (line 11-12) and stores them as pairs in the buffer $\mathcal{D_{\tau}}$ (line 13).
(2) Weight Sharing (line 14-16)
We sort the candidate pairs in $\mathcal{D_{\tau}}$ in descending order of their matching scores $s_{m}$ (line 14). Then, the top-N pairs $\mathcal{N}$ are selected according to the desired sharing ratio $\alpha$ (line 15). Finally, for each selected head pair, we keep the weight values of the head in former layer unchanged and select to change the weight values of the head in latter layer (line 16), resulting in the final model $\mathcal{M^{*}}$.
Input: top-N matched head pairs $\mathcal{N}$, Original LLM $\mathcal{M}$
Output: The LLM $\mathcal{M^{*}}$ after weight sharing
1 Initialize one dict $\mathcal{D}$;
2 for pair in $\mathcal{N}$ do
3      $index_{i}$, $index_{j}$ $\leftarrow pair$;
4      $(layer_{i},i)\leftarrow index_{i}$;
5      $(layer_{j},j)\leftarrow index_{j}$;
6      if $layer_{i}<layer_{j}$ then
7           $\mathcal{D}[index_{i}]=index_{j}$;
8
9     else
10           $\mathcal{D}[index_{j}]=index_{i}$;
11
12
13Sort $\mathcal{D}$ by the key;
14 $\mathcal{M^{*}}\leftarrow\mathcal{M}$;
15 for $head_{i},head_{j}$ in $\mathcal{D}$ do
16      $\mathcal{M^{*}}\leftarrow$ $\mathcal{M^{*}}$, $head_{i}$, $head_{j}$);
17
18return $\mathcal{M}^{*}$;
Algorithm 2 Weight_Share Function
Appendix B Relation between Matrix Weights and Attention Map Similarities
Our visual analysis (in Section3.2) can intuitively illustrate that the shared head pair with the highest similar weight matrices to a large degree exhibits highly similar attention maps across different datasets. Aside from this, we also provide a metric to measure the relation:
$Degree=\frac{\#num~{}of~{}matched~{}heads}{\#num~{}of~{}samples}$ (6)
Each time we use 100 randomly selected samples from each dataset to calculate the average cosine similarity value and repeat 5 times. According to the sharing ratio $\alpha$, we will respectively select the top-K shared head pairs based on the weight matrix similarity and attention map similarity (calculated cosine similarity as metric), denoted as Set $A$ and $B$. The number of matched heads is calculated as the cardinality of the intersection of Set $A$ and Set $B$, denoted as $|A\cap B|$. From Table 6, it can be seen the matched degree is relatively high, so it is reasonable for us to do weight sharing directly across heads.
Dataset WMT-14 CQA WSC Average
Llama2-7B 100% 91.30% 100% 97.10%
Table 6: The degree metric measured to quantify the relation between matrix weights and attention map similarities. CQA stands for CommonsenseQA benchmark.
Appendix C Implementation Details
@@
C.1 Baselines
To our knowledge, there is no existing baseline for our methods, due to the absence of prior research on fine-grained weight sharing for LLMs. To provide a comprehensive demonstration of the effectiveness of our DirectShare, we can choose another important memory-efficient method of a different category for comparison. Here, we select two model pruning methods applied in LLMs: one classical model pruning method Magnitude Pruning and one state-of-the-art structured pruning method LLM-Pruner. We do not consider unstructured pruning methods in this paper since they can not achieve real memory reduction without specialized hardware or software. Appendix G lists more comparison results with some baselines modified to support LLMs.
@@ C.1 Baselines
C.2 Benchmarks
Logical and Common Sense Reasoning. In the domain of reasoning, we consider two Chinese natural language inference benchmarks and three English benchmarks: CMNLI (Xu et al., 2020), OCNLI (Hu et al., 2020), along with AX-b, AX-g and RTE from SuperGLUE (Wang et al., 2019).
Knowledge-related Tasks. We perform evaluations regarding knowledge on various datasets: WinoGrande (Levesque et al., 2012) about language, BoolQ (Clark et al., 2019) testing knowledge question answering, C-Eval (Huang et al., 2023) and MMLU (Hendrycks et al., 2021) standing for two comprehensive examination benchmarks.
C.3 Post-training Details
For carrying out the post-training process, we employ the code framework from LLaMA-Factory repository [^2] https://github.com/hiyouga/LLaMA-Factory with DeepSpeed ZeRO-1333Because of our designed special loss function in the post-training stage, only DeepSpeed ZeRO-1 can work.. The Adam optimizer with a learning rate of 5e-5 is used in our experiment and the parameter values assigned during training are $\beta_{1}=0.9$ and $\beta_{2}=0.95$. For Llama 2-7B model, we set the batch size to 32. While for Llama 2-13B model, the batch size of training is only 8 subject to the limited computational resources. Besides, the maximum context size and $\gamma$ are set to 4096 and 0.5, respectively.
Appendix D Experimental Results based on Baichuan 2 Models
@@
Table 7: Evaluation results on reasoning tasks when applying DirectShare to Baichuan2 models.
@@
Table 8: NLU abilities of Baichuan2 models after DirectShare.
D.1 Logical and Common Sense Reasoning
Table 7 presents a comparison on five datasets about reasoning abilities for three memory-efficient methods performed on the Baichuan2 models. Our results show that compared to NLU and knowledge-related abilities (listed in Table 8,10), DirectShare can indeed maintain its reasoning abilities to a large extent. Specifically, at 30% ratio, DirectShare remains competitive with LLM-Pruner.
D.2 Natural Language Understanding
0% Llama 2-13B 55.44 71.50 40.17 55.81 60.24 58.03 42.40 60.00
30% DirectShare 50.18 59.36 22.30 30.79 26.53 27.53 27.40 27.80
PostShare 53.68 $\uparrow 3.50$ 71.25 $\uparrow 11.89$ 25.80 $\uparrow 3.50$ 33.90 $\uparrow 3.11$ 32.03 $\uparrow 3.30$ 29.07 $\uparrow 1.54$ 33.60 $\uparrow 6.20$ 38.80 $\uparrow 11.00$
Table 9: Performance of PostShare based on the Llama2-13B backbone. * means choosing relatively good performance across different training steps.
@@ D.2 Natural Language Understanding
Table 10: Results on knowledge-related tasks of Baichuan2 models after DirectShare.
D.3 Knowledge-related Tasks
The results of Baichuan2 models on knowledge-related tasks are shown in Table 10. Similar decline appears in Llama2-7B/13B models as well.
Appendix E PostShare on Llama 2-13B Model
@@
Appendix F More Analysis
F.1 Training Time Costs of PostShare
For Llama2-7B model, our full-parameter post-training in English Wikipedia corpus uses around 11.27 hours for 0.5 epoch due to our limited computational resources. We list the post-training experimental settings in Appendix C.3. In the same settings, the original full-parameter post-training requires 10.78 hours for 0.5 epoch, which indicates our PostShare does not increase much training costs ($\approx$30min).
F.2 Overfitting Phenomenon in PostShare
@@
CMNLI OCNLI AX-b AX-g RTE RACE- middle RACE- high OBQA CSL TNEWS Wino- Grande BoolQ C-Eval MMLU Avg.
Choosing A 33.33 32.50 57.07 51.68 49.10 21.45 21.53 26.00 51.25 20.22 50.18 54.43 26.24 26.53 37.25
Choosing B 33.59 31.25 51.09 53.09 53.43 25.00 25.53 22.20 49.38 12.20 52.63 57.22 28.27 25.63 37.18
Table 13: Performance comparison between choosing different shared weights for a pair of matched heads A, B.
(a) Sharing Ratio=20%
(b) Sharing Ratio=30%
Figure 7: Ratios of weight sharing across the MHA Layers in Llama2-7B/13B and Baichuan2-7B/13B.
F.3 Impact of Different Head-wise Matching Functions
F.4 How to Choose the Shared Weights?
We select the weight value from the head in the former layer as the shared weights. The reason behind is that we observe the weight values of the earlier layers are more sensitive in our experiments, which is also supported by previous studies (Men et al., 2024; Chen et al., 2024). Thus, we tend to keep the weight values of the head in former layer unchanged. We also additionally make a performance comparison between choosing different shared weights for a pair of matched heads A, B. Suppose head A is in the former layer and head B is in the latter layer. In Table 13, it can be seen that the performance gap is small on the downstream tasks whether choosing A or B, but the overall performance is a little better when choosing the head weight in the former layers as the shared weights.
F.5 Latency Evaluations
Our paper aims to reduce the memory requirements for LLMs instead of accelerating. Since weight sharing does not change the total matrix multiplications, DirectShare can theoretically maintain the original model inference speed but our current implementation of head-wise weight sharing is hard to save memory and speed up at the same time, which may be a programming skill issue. Table 16 reports the inference speed tested under the test set of WikiText2. We will consider further optimization in future work to achieve a balance between memory and speedup.
Method CMNLI OCNLI AX-b AX-g RTE RACE- middle RACE- high OBQA CSL TNEWS Wino- Grande BoolQ C-Eval MMLU
DirectShare Ratio=10% 33.00 32.50 54.17 51.97 50.90 28.34 28.96 28.20 54.37 20.86 52.63 67.74 28.75 43.43
GQA group size=2 32.20 27.50 58.33 50.00 47.29 25.21 25.61 22.20 46.25 6.92 50.53 37.83 25.68 25.80
DirectShare Ratio=30% 33.33 32.50 57.07 51.68 49.10 21.45 21.53 26.00 51.25 20.22 50.18 54.43 26.24 26.53
Table 14: Comparison results with GQA across various datasets.
Ratio CMNLI OCNLI AX-b AX-g RTE RACE- middle RACE- high OBQA CSL TNEWS Wino- Grande BoolQ C-Eval MMLU
Llama2-7B 0% 32.98 33.12 53.53 55.34 49.82 33.15 35.51 31.80 55.62 20.22 54.04 70.67 32.20 46.69
DirectShare 10% 33.00 32.50 54.17 51.97 50.90 28.34 28.96 28.20 54.37 20.86 52.63 67.74 28.75 43.43
SEQ 10% 31.55 37.50 49.82 51.40 52.35 28.62 28.39 27.00 57.50 15.39 51.23 64.77 30.40 39.69
DirectShare 30% 33.33 32.50 57.07 51.68 49.10 21.45 21.53 26.00 51.25 20.22 50.18 54.43 26.24 26.53
SEQ 30% 31.72 32.50 47.55 50.00 46.93 22.63 22.30 26.20 50.62 9.83 49.47 52.60 22.05 23.49
Table 15: Comparison results with SEQ across various datasets.
Model Sharing Ratio Speed
Llama2-7B 0% 41.70 tokens/s
30% MHA 8.52 tokens/s
30% MHA+FFN 8.34 tokens/s
Llama2-13B 0% 24.20 tokens/s
30% MHA 5.63 tokens/s
30% MHA+FFN 5.67 tokens/s
Table 16: The inference speed of our method tested on a single A800-80GB GPU.
F.6 Visualization Study on the Shared Weights